OpenStudy (anonymous):

The tip of a 15-inch wiper blade wipes a path that is 36 inches long. What is the angle of rotation of the blade in radians to the nearest tenth? 2.4 radians 1.2 radians 2.8 radians 0.4 radians

OpenStudy (dangerousjesse):

\(\textsf{s = length of arc, which is = 36}\) \(\textsf{rotation (r) = 15}\) so \(\textsf{s=rθ}\) Say \(\textsf{36=15θ}\) \(θ=\frac{36}{15}\)

OpenStudy (dangerousjesse):

Refresh the page if you see "��" :)

OpenStudy (anonymous):

I don't get it

OpenStudy (phi):

There are (at least) 2 ways to do this. one way is to *memorize* the formula arc length s = radius * angle_in_radians written as \[ s = r \ \theta\] for your problem, "solve for θ" \[ \theta= \frac{s}{r} \] and remember this *only works for theta in radians* (not degrees)

OpenStudy (phi):

the other way is to use ratios. we know all the way around the circle takes 2 pi radians (i.e. 360 degrees) the circumference (all the way round) is 2 pi r now write a ratio of angle/distance = angle/distance where the first ratio is the whole circle: \[ \frac{2 \pi}{2 \pi r} = \frac{\theta}{s} \] the left side simplifies to \[ \frac{1}{r} = \frac{\theta}{s} \] if we cross multiply, we get \[ s = r \ \theta\] which seems vaguely familiar

OpenStudy (phi):

using a radius r=15 inches, all the way round your circle i.e. its circumference is 2 pi * 15 = 30 pi inches (about 94.25 inches) 36 inches will only be part way round as a ratio: \[ \frac{2 \pi \ radians}{30 \pi \ inches} = \frac{\theta}{36 \ inches} \] or, after simplifying \[ \frac{1 \ radian}{15} = \frac{\theta}{36 } \] multiply both sides by 36 to get \[ \frac{36}{15} \ radians= \theta \] so theta is 12/5 radians. as a decimal, it is 2.4 radians